The problem is talking about surface area and the number for height is 2ft and for length is 3ft and the width is 3ft and the answer i got is 48ft squared
Answer by fcabanski(1391)

The surface area of a rectangular prism (a 3D rectangular box) is 2ab + 2ac + 2bc.


In this problem call a = 2, b=3, c=3.
The surface area is 2*2*3 + 2*2*3 + 2*3*3 = 12 + 12 + 18 = 42 square feet.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com